iet x = 2 in (* x 2) in let x Select one: O a. 2 O b. 6 O d. 4
Answer #1

initial value of x is 2

then x = x * 2 = 2 * 2 = 4

An suer

4

\color{red}Note: \;Please \;comment \;below \;if \;you \;have \;any \;doubts